You've answered the question regarding the probability of a yarborough in bridge. What is the probability that the highest ranking card is a 7?
Answer by stanbon(75887)

a yarborough is a 13-card hand containing only cards numbered 2 through 9.
---------------
What is the probability that the highest ranking card is a 7?
---
# of cards that are 2 through 7 = 6*4 = 24
# of cards that are 2 through 9 = 8*4 = 32
----
# of yarborough hands = 32C13
# of yarborough hands with max card = 7 is 24C13
Assuming you have a yarborough the answer is 24C13/32C13 = 0.007
---
Assuming no conditions the prob of having yarboroug
with max card = 7 is 24C13/52C13 = 0.000003931...
Cheers,
Stan H.